OpenStudy (anonymous):

Convert from Cartesian to Cylindrical coordates in the following triple integral: 1 / 0, (sqrt(1-x^2)) / -(sqrt(1-x^2)), 3/0 of (x + y^2)/(z^2 +1) dzdydx

OpenStudy (anonymous):

\[\int\limits_{1}^{0}\int\limits_{-\sqrt{1-x^2}}^{\sqrt{1-x^2}}\int\limits_{0}^{3} (x+y^2)/(z^2 +1) dzdydx\]

OpenStudy (anonymous):

err that dx should be from 1 to 0

OpenStudy (anonymous):

I think it's 0 <= r <= 1; 0 <= theta <= pi/2; 0 <= z <= 3